Checking my work-3 to the 8th power is 6561? 7 to the 5th =2401?
Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(30993)

Checking my work-3 to the 8th power is 6561? 7 to the 5th =2401
-----------------
2401 is 7^4
7^5 = 16,807

Answer by checkley77(12569)

-3^8=6,561
7^5=16,807